OpenStudy (anonymous):

If f(x)=(x+1)^−1 and g(x)=x−2, what is the domain of f(x)÷g(x)?

OpenStudy (anonymous):

all values of x (−∞,−1),(−1,2), and (2,∞) (−∞,2) and (2,∞) (−∞,−1] and [2,∞)

OpenStudy (anonymous):

first off, can you write the function f(x) ÷ g(x)?

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

\[f(x) \div g(x)=(x+1)^{-1}\div (x-2)=\frac{ 1 }{ x+1 }\div (x-2)\]

OpenStudy (anonymous):

do the division to right that as a single fraction.

OpenStudy (anonymous):

i have no clue how to do any this

OpenStudy (anonymous):

do you know how to divide fractions? because that's what this is...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you need to flip the (x-2) part to it's reciprocal and change the division sign to multiplication, then just multiply across like these examples.